# An Information-Theoretic Framework for Receiver Quantization in Communication

### v2

@@ III Achievable Rate Under Receiver Quantization: Exact and Asymptotic Results
From Theorem 2, we immediately obtain the following corollary, which shows that the parameter $\gamma$ dominates the asymptotic behavior of performance, especially the low-SNR slope of achievable rate and the saturation rate.
@@
    So the saturation rate (i.e. the high-SNR limit of achievable rate) is given by
@@ IV-A Achievable Rate Evaluation under Uniform Quantization
Clearly, decreasing $L$ reduces the granular distortion but increases the overload distortion. When $L<L^{*}$ the overload distortion increases quickly, thereby causing the waterfall of achievable rate.
@@ IV-B Consistency Between Rate Maximization and MSE Minimization and a GMI-MMSE Formula
Lemma 7: For uniform quantization at the receiver, for $b>1$ (i.e., except for one-bit quantization), we have
@@
Theorem 8: For the channel (4) where $q(\cdot)$ is the uniform quantizer described in Sec. II-A2 with resolution $b>1$, the input employs i.i.d. complex Gaussian codebook, and the receiver employs nearest neighbor decoding rule (7), we have the following properties.
@@
    which is due to suboptimal loading factor and can be reduced by improving the accuracy of gain control. Numerical results in Fig. 4 and Fig. 5 show the importance of reducing loading loss.
@@
In Table I, [^8] Some numerical results therein have been given in [References, Table II], in which an error occurred in actual SNR computation for Gaussian source, $b=8$. we show numerical results of optimal uniform quantization under our transceiver architecture (an approximation of $\mathsf{mmse}$ given by $\ln(2K)/3K^{2}$, and an approximation of $\bar{I}_{\textrm{GMI}}$, denoted by $\hat{I}_{\textrm{GMI}}$, will be introduced in Sec. VI-B).
@@ IV-C Geometry of Optimal Uniform Quantization at the Receiver
Now we are ready to illustrate the geometry of optimal uniform quantization at the receiver in the $N$-dimensional Euclidean space, as shown in Fig. 7. We use boldface letters to denote codewords or signal vectors in the equivalent model (69), e.g., $\mathbf{X}=[\mathsf{X}_{1},...,\mathsf{X}_{N}]$. Since i.i.d. codebook is considered, the vector $\mathbf{X}$ and other vectors in Fig. 7 are all i.i.d. random vectors. Thus, we have $\mathrm{E}\left[\|\mathbf{X}\|^{2}\right]=N\mathcal{E}_{x}$, and the empirical average power of the input codeword converges in probability to $\mathcal{E}_{x}$, i.e.,
@@
Thus, the lengths of $\mathbf{V}$, $\mathbf{W}$, and $\mathbf{Y}$ are $\sqrt{N(\mathcal{E}_{x}+\mathcal{\sigma}^{2})}$, $\sqrt{N\cdot\mathsf{MMSE}}$, and $\sqrt{N(\mathcal{E}_{x}+\mathcal{\sigma}^{2}-\mathsf{MMSE})}$, respectively. We then have two right triangles in Fig. 7: the quantization triangle $\mathbf{0VY}$ [^9] Note that the angle between $\mathbf{Y}$ and $\mathbf{V}-\mathbf{Y}$ is a right angle. Since Fig. 7 is a one-dimensional illustration of the geometry in the $N$-dimensional space, we cannot make every right angle therein look like a right angle., which determines the saturation rate $\bar{I}^{*}_{\textrm{GMI}}$ in (65), and the noise triangle $\mathbf{0XV}$, which determines the channel capacity $C$, i.e., the limit of achievable rate under fine quantization. These two triangles jointly determine the triangle $\mathbf{0XY}$, and consequently determine the achievable rate $I^{*}_{\textrm{GMI}}$.
We note that $\mathbf{YVX}$ and $\mathbf{0XY}$ are not right triangles. In fact, for finite $\mathsf{SNR}$ and finite resolution, $\mathrm{E}\left[|\mathsf{Y}-\mathsf{X}|^{2}\right]$ is strictly smaller than $\mathrm{E}\left[|\mathsf{Z}|^{2}\right]+\mathrm{E}\left[|\mathsf{W}|^{2}\right]$, because the error $\mathsf{W}$ is correlated with the noise $\mathsf{Z}$. Treating the error as independent noise always reduces the achievable rate.
@@
The high-resolution limit of achievable rate (i.e., the capacity) can be expressed as
@@
The high-SNR limit of achievable rate (i.e., the saturation rate) can be expressed as
@@ V Asymptotic Analysis: Characterization of $I_{\textrm{GMI}}(L)$
1) Overload Region ($L<L^{*}$): In this region overload distortion dominates the performance and the achievable rate behaves like a waterfall. In particular, for high-resolution quantization at high SNR, the GMI is approximated by $\log(L^{3}/4\phi(L))$ (except when $L$ is very close to $L^{*}$), which appears almost linear; see (119a) and Fig. 5, (d)-(f). When $L$ approaches $L^{*}$, the impact of granular distortion becomes increasingly important, so that the GMI grows slower and approaches its maximum. As $L\to 0$, the effective resolution of the quantizer reduces to one bit and $I_{\textrm{GMI}}(L)$ converges to $I_{\textrm{GMI}}^{\textrm{1-bit}}$ given in (35).
2) Underload Region ($L>L^{*}$): In this region granular distortion dominates the performance. As $L$ increases, the rate loss increases like $\log(1+L^{2}K^{-2}\mathsf{SNR}/12)$ (see (120)) over a wide range. When $L\gg L^{*}$, the effective resolution reduces to one bit and $I_{\textrm{GMI}}(L)$ converges to $I_{\textrm{GMI}}^{\textrm{1-bit}}$ again.
@@ V-A Decay of Rate Loss under High-Resolution Receiver Quantization
The following result shows that, the rate loss due to only overload distortion decays exponentially as the loading factor increases. Asymptotically, the loading loss is directly proportional to the overload distortion, and is also directly proportional to the SNR.
@@
In the high-resolution limit, granular distortion vanishes and the normalized MSE includes only the overload distortion, namely $\underline{\mathsf{mse}}_{\textrm{o}}$. From Proposition 5, (45c), and (46c), we obtain another expression of the overload distortion as
@@
On the other hand, the rate loss due to only granular distortion decays quadratically as the step size vanishes. To prove this we need the following lemma which is one of the various forms of the Euler-Maclaurin summation formula.
@@
in which the first two terms of the RHS come from (94a), and the last term comes from (105). As shown in Fig. 9, the proposed formula well approximates the transition from overload region to underload region.
@@ V-B2 Underload Region
which is highly accurate even when the resolution is low; see Fig. 11 and Fig. 11. We have also considered an even simpler approximation $C-L^{2}K^{-2}\mathsf{SNR}/12\;\textrm{nats/c.u.}$ in the high-resolution scenario; see Fig. 9 and Fig. 9. However, it becomes less useful as the loading loss increases, as shown in Fig. 11 and Fig. 11. The approximation (120) also implies that the saturation rate can be approximated by
@@ VI-A Approximations of Optimal Loading Factor
We first provide a new proof of the property (60) for the step size and loading factor of the optimal uniform quantizer. In [68], it was shown that the uniform quantizer that minimizes the MSE must satisfy (60) if the input density has infinite support. Here we show that (60) is a natural corollary of the achievable rate results in Theorem 12 and Theorem 14.
@@ VII-A Relationship Between MSE and Achievable Rate: Numerical Examples
Remark: In achievable rate evaluation, if we use the additive noise model of the quantizer, i.e., treating the MSE as additive Gaussian noise, then we obtain an estimate as
@@ VII-C On Possible Gain of Further Quantization Rules
where $\mathsf{mmse}^{\textrm{nu}}$ is the normalized minimum MSE of all possible nonuniform quantizers. This result can also be understood via a similar geometrical interpretation as that given in Sec. IV-C. From (154) we can infer that, under our setting of transceiver architecture, the maximum gain in achievable rate from replacing the optimal uniform quantizer by the optimal nonuniform quantizer is determined by the reduction of the normalized MMSE from $\mathsf{mmse}$ to $\mathsf{mmse}^{\textrm{nu}}$. The numerical results in Fig. 17(a) and Fig. 17(b) show that the rate gain is marginal for $b=2$. For higher resolutions the rate gain grows slowly. When $b=4$, according to the numerical results in [31], the optimal non-uniform quantization improves the saturation rate $\log(1/\gamma)$ by no more than $4.4\%$, and optimized quantization with equispaced thresholds improves the saturation rate by $1.5\%$. In fact, for nonuniform quantization with Gaussian input, the asymptotic quantization theory, in particular the Panter-Dite formula [74], yields a “6-dB-per-bit” rule [^13] There is another version of the $6$-dB-per-bit rule as $B_{\mathrm{eff}}=(\mathsf{SNR}_{q}-1.76)/6.02$, which has been widely used in practice to calculate the effective number of bits (ENoB) of a quantizer [2, 3]. It is derived under assumptions that the granular distortion is uniformly distributed over $\ell/2,\ell/2]$ and there is no overload distortion; see [2] for more details. for the maximum SNR as [42]
@@
where $\mathsf{SNR}_{q}^{\textrm{nu}}=1/\mathsf{mmse}^{\textrm{nu}}$. Comparing with (141), we can infer that under our setting the rate gain is limited in the range of resolution of interest. In particular, the gain in maximum saturation rate (65) is $\log\left(\mathsf{SNR}_{q}^{\textrm{nu}}/\mathsf{SNR}_{q}\right)$, which increases logarithmically with $b$.
@@
A final remark is about dithering, an intentional randomization technique in source quantization [42], which may also be beneficial in receiver quantization; see, e.g., [32], which showed that nonsubtractive Gaussian dithering [^16] In nonsubtractive dithering, one adds a dither signal $W_{\mathrm{d}}$ independent of the quantizer input $V$ before quantization, yielding an output $q(V+W_{\mathrm{d}})$. Subtractive dithering has an additional step that subtracts the dither signal after quantization and finally obtains $q(V+W_{\mathrm{d}})-W_{\mathrm{d}}$. See [42] for more discussion. improves GMI under certain circumstances. In our setting, although nonsubstractive dithering reduces the SNR, it also changes $\gamma$ by enlarging the standard deviation of the input so that all normalized thresholds and levels are scaled by a factor smaller than one. The overall impact on GMI is thus unclear because it depends on the change of the product $\gamma\cdot\mathsf{SNR}$. Nevertheless, for uniform quantization, if we employ perfect gain control which guarantees $\gamma\equiv\mathsf{mmse}$, then nonsubstractive Gaussian dithering is always harmful since the GMI (54) satisfies
@@ VIII Concluding Remarks
The goal of this study is to evaluate the impact of resolution reduction on information-theoretic limits of communications with receiver quantization. Leveraging the GMI as a basic tool, which enables us to take the decoding rule into consideration, we establish an array of exact and asymptotic results under a standard transceiver architecture. Our results indicate a critical issue in system design that arises as the resolution decreases, namely, optimizing the loading factor by gain control, which minimizes the loss in achievable rate by eliminating the loading loss. The remaining irreducible loss is an appropriate evaluation of the inherent robustness of the considered transceiver architecture. Our results also establish explicit connections between the MSE (affected by the gain control) and the rate loss. Although for general receiver quantization rules, smaller MSE does not necessarily imply smaller rate loss, for the commonly used uniform quantizer we prove that the unique loading factor that minimizes the MSE also maximizes the GMI (i.e., achieves zero loading loss). For perfect gain control, we show that the irreducible loss is determined by the product of the normalized MMSE and the SNR, and provide a geometrical interpretation for this result. Performance approximations and per-bit rules in this case are also given. For imperfect gain control, to understand its impact, we characterize the decay of small rate loss in the high-resolution regime, and propose approximations of the achievable rate as a function of the loading factor which are fairly accurate for moderate resolutions. These results provide insight into transceiver design with nonnegligible quantization effect, especially regarding the choice of quantizer resolution and the design of AGC.
A limitation of this work is that the obtained analytical results are derived under the Gaussian codebook. Like many classical information-theoretic results given by mutual information, for the GMI, it is also not easy to get closed-form or analytical results without assuming Gaussian input. For practical systems with finite alphabet, we can expect that our result applies well for Gaussian-like constellations that come from constellation shaping. But it is natural to ask whether our results are still useful in the cases of QAM, phase-shift keying (PSK), and other commonly used inputs, which are all bounded (unlike the unbounded Gaussian input). Since in these cases Proposition 1 does not apply, the GMI evaluation may rely heavily on numerical computation. We left this problem to future study. Nevertheless, the insight gained in this work will be very helpful for the finite-alphabet case.

### v3

@@ III Achievable Rate Under Receiver Quantization: Exact and Asymptotic Results
From Theorem 2, we immediately obtain the following corollary, which shows that the parameter $\gamma$ dominates the asymptotic behavior of performance, especially the low-SNR slope of the achievable rate and the saturation rate.
@@
    So the saturation rate (i.e. the high-SNR limit of the achievable rate) is given by
@@ IV-A Achievable Rate Evaluation under Uniform Quantization
Clearly, decreasing $L$ reduces the granular distortion but increases the overload distortion. When $L<L^{*}$ the overload distortion increases quickly, thereby causing the waterfall of the achievable rate.
@@ IV-B Consistency Between Rate Maximization and MSE Minimization and a GMI-MMSE Formula
Lemma 7: For uniform quantization at the receiver, for $b>1$ (i.e., except for the one-bit quantization), we have
@@
Theorem 8: For the channel (4) where $q(\cdot)$ is the uniform quantizer described in Sec. II-A2 with resolution $b>1$, the input employs i.i.d. complex Gaussian codebook, and the receiver employs the nearest neighbor decoding rule (7), we have the following properties.
@@
    which is due to a suboptimal loading factor and can be reduced by improving the accuracy of gain control. Numerical results in Fig. 4 and Fig. 5 show the importance of reducing loading loss.
@@
In Table I, [^8] Some numerical results therein have been given in [References, Table II], in which an error occurred in actual SNR computation for Gaussian source, $b=8$. we show numerical results of the optimal uniform quantization under our transceiver architecture (an approximation of $\mathsf{mmse}$ given by $\ln(2K)/3K^{2}$, and an approximation of $\bar{I}_{\textrm{GMI}}$, denoted by $\hat{I}_{\textrm{GMI}}$, will be introduced in Sec. VI-B).
@@ IV-C Geometry of Optimal Uniform Quantization at the Receiver
Now we are ready to illustrate the geometry of the optimal uniform quantization at the receiver in the $N$-dimensional Euclidean space, as shown in Fig. 7. We use boldface letters to denote codewords or signal vectors in the equivalent model (69), e.g., $\mathbf{X}=[\mathsf{X}_{1},...,\mathsf{X}_{N}]$. Since i.i.d. codebook is considered, the vector $\mathbf{X}$ and other vectors in Fig. 7 are all i.i.d. random vectors. Thus, we have $\mathrm{E}\left[\|\mathbf{X}\|^{2}\right]=N\mathcal{E}_{x}$, and the empirical average power of the input codeword converges in probability to $\mathcal{E}_{x}$, i.e.,
@@
Thus, the lengths of $\mathbf{V}$, $\mathbf{W}$, and $\mathbf{Y}$ are $\sqrt{N(\mathcal{E}_{x}+\mathcal{\sigma}^{2})}$, $\sqrt{N\cdot\mathsf{MMSE}}$, and $\sqrt{N(\mathcal{E}_{x}+\mathcal{\sigma}^{2}-\mathsf{MMSE})}$, respectively. We then have two right triangles in Fig. 7: the quantization triangle $\mathbf{0VY}$ [^9] Note that the angle between $\mathbf{Y}$ and $\mathbf{V}-\mathbf{Y}$ is a right angle. Since Fig. 7 is a one-dimensional illustration of the geometry in the $N$-dimensional space, we cannot make every right angle therein look like a right angle., which determines the saturation rate $\bar{I}^{*}_{\textrm{GMI}}$ in (65), and the noise triangle $\mathbf{0XV}$, which determines the channel capacity $C$, i.e., the limit of the achievable rate under fine quantization. These two triangles jointly determine the triangle $\mathbf{0XY}$, and consequently determine the achievable rate $I^{*}_{\textrm{GMI}}$.
We note that $\mathbf{YVX}$ and $\mathbf{0XY}$ are not right triangles. In fact, for finite $\mathsf{SNR}$ and finite resolution, $\mathrm{E}\left[|\mathsf{Y}-\mathsf{X}|^{2}\right]$ is strictly smaller than $\mathrm{E}\left[|\mathsf{Z}|^{2}\right]+\mathrm{E}\left[|\mathsf{W}|^{2}\right]$, because the error $\mathsf{W}$ is correlated with the noise $\mathsf{Z}$. Treating the error as an independent noise always reduces the achievable rate.
@@
The high-resolution limit of the achievable rate (i.e., the capacity) can be expressed as
@@
The high-SNR limit of the achievable rate (i.e., the saturation rate) can be expressed as
@@ V Asymptotic Analysis: Characterization of $I_{\textrm{GMI}}(L)$
1) Overload Region ($L<L^{*}$): In this region the overload distortion dominates the performance and the achievable rate behaves like a waterfall. In particular, for high-resolution quantization at high SNR, the GMI is approximated by $\log(L^{3}/4\phi(L))$ (except when $L$ is very close to $L^{*}$), which appears almost linear; see (119a) and Fig. 5, (d)-(f). When $L$ approaches $L^{*}$, the impact of granular distortion becomes increasingly important, so that the GMI grows slower and approaches its maximum. As $L\to 0$, the effective resolution of the quantizer reduces to one bit and $I_{\textrm{GMI}}(L)$ converges to $I_{\textrm{GMI}}^{\textrm{1-bit}}$ given in (35).
2) Underload Region ($L>L^{*}$): In this region the granular distortion dominates the performance. As $L$ increases, the rate loss increases like $\log(1+L^{2}K^{-2}\mathsf{SNR}/12)$ (see (120)) over a wide range. When $L\gg L^{*}$, the effective resolution reduces to one bit and $I_{\textrm{GMI}}(L)$ converges to $I_{\textrm{GMI}}^{\textrm{1-bit}}$ again.
@@ V-A Decay of Rate Loss under High-Resolution Receiver Quantization
The following result shows that, the rate loss due to only the overload distortion decays exponentially as the loading factor increases. Asymptotically, the loading loss is directly proportional to the overload distortion, and is also directly proportional to the SNR.
@@
In the high-resolution limit, the granular distortion vanishes and the normalized MSE includes only the overload distortion, namely $\underline{\mathsf{mse}}_{\textrm{o}}$. From Proposition 5, (45c), and (46c), we obtain another expression of the overload distortion as
@@
On the other hand, the rate loss due to only the granular distortion decays quadratically as the step size vanishes. To prove this we need the following lemma which is one of the various forms of the Euler-Maclaurin summation formula.
@@
in which the first two terms of the RHS come from (94a), and the last term comes from (105). As shown in Fig. 9, the proposed formula well approximates the transition from the overload region to the underload region.
@@ V-B2 Underload Region
which is highly accurate even when the resolution is low; see Fig. 11 and Fig. 11. We have also considered an even simpler approximation $C-L^{2}\mathsf{SNR}/12K^{2}\;\textrm{nats/c.u.}$ in the high-resolution scenario; see Fig. 9 and Fig. 9. However, it becomes less useful as the loading loss increases, as shown in Fig. 11 and Fig. 11. The approximation (120) also implies that the saturation rate can be approximated by
@@ VI-A Approximations of Optimal Loading Factor
We first provide a new proof of the property (60) for the step size and loading factor of the optimal uniform quantizer. In [68], it was shown that the uniform quantizer that minimizes the MSE must satisfy (60) if the input density has an infinite support. Here we show that (60) is a natural corollary of the achievable rate results in Theorem 12 and Theorem 14.
@@ VII-A Relationship Between MSE and Achievable Rate: Numerical Examples
Remark: In the achievable rate evaluation, if we use the additive noise model of the quantizer, i.e., treating the MSE as additive Gaussian noise, then we obtain an estimate as
@@ VII-C On Possible Gain of Further Quantization Rules
where $\mathsf{mmse}^{\textrm{nu}}$ is the normalized minimum MSE of all possible nonuniform quantizers. This result can also be understood via a similar geometrical interpretation as that given in Sec. IV-C. From (154) we can infer that, under our setting of transceiver architecture, the maximum gain in achievable rate from replacing the optimal uniform quantizer by the optimal nonuniform quantizer is determined by the reduction of the normalized MMSE from $\mathsf{mmse}$ to $\mathsf{mmse}^{\textrm{nu}}$. The numerical results in Fig. 17(a) and Fig. 17(b) show that the rate gain is marginal for $b=2$. For higher resolutions the rate gain grows slowly. When $b=4$, according to the numerical results in [31], the optimal non-uniform quantization improves the saturation rate $\log(1/\gamma)$ by no more than $4.4\%$, and optimized quantization with equispaced thresholds improves the saturation rate by $1.5\%$. In fact, for nonuniform quantization with a Gaussian input, the asymptotic quantization theory, in particular the Panter-Dite formula [74], yields a “6-dB-per-bit” rule [^13] There is another version of the $6$-dB-per-bit rule as $B_{\mathrm{eff}}=(\mathsf{SNR}_{q}-1.76)/6.02$, which has been widely used in practice to calculate the effective number of bits (ENoB) of a quantizer [2, 3]. It is derived under assumptions that the granular distortion is uniformly distributed over $\ell/2,\ell/2]$ and there is no overload distortion; see [2] for more details. for the maximum SNR as [42]
@@
where $\mathsf{SNR}_{q}^{\textrm{nu}}=1/\mathsf{mmse}^{\textrm{nu}}$. Comparing with (141), we can infer that under our setting the rate gain is limited in the range of resolution of interest. In particular, the gain in the maximum saturation rate (65) is $\log\left(\mathsf{SNR}_{q}^{\textrm{nu}}/\mathsf{SNR}_{q}\right)$, which increases logarithmically with $b$.
@@
A final remark is about dithering, an intentional randomization technique in source quantization [42], which may also be beneficial in receiver quantization; see, e.g., [32], which showed that nonsubtractive Gaussian dithering [^16] In nonsubtractive dithering, one adds a dither signal $W_{\mathrm{d}}$ independent of the quantizer input $V$ before quantization, yielding an output $q(V+W_{\mathrm{d}})$. Subtractive dithering has an additional step that subtracts the dither signal after quantization and finally obtains $q(V+W_{\mathrm{d}})-W_{\mathrm{d}}$. See [42] for more discussion. improves the GMI under certain circumstances. In our setting, although nonsubstractive dithering reduces the SNR, it also changes $\gamma$ by enlarging the standard deviation of the input so that all normalized thresholds and levels are scaled by a factor smaller than one. The overall impact on GMI is thus unclear because it depends on the change of the product $\gamma\cdot\mathsf{SNR}$. Nevertheless, for uniform quantization, if we employ perfect gain control which guarantees $\gamma\equiv\mathsf{mmse}$, then nonsubstractive Gaussian dithering is always harmful since the GMI (54) satisfies
@@ VIII Concluding Remarks
The goal of this study is to evaluate the impact of resolution reduction on information-theoretic limits of communications with receiver quantization. Leveraging the GMI as a basic tool, which enables us to take the decoding rule into consideration, we establish an array of exact and asymptotic results under a standard transceiver architecture. Our results indicate a critical issue in system design that arises as the resolution decreases, namely, optimizing the loading factor by gain control, which minimizes the loss in achievable rate by eliminating the loading loss. The remaining irreducible loss is an appropriate evaluation of the inherent robustness of the considered transceiver architecture. Our results also establish explicit connections between the MSE (affected by the gain control) and the rate loss. Although for general receiver quantization rules, a smaller MSE does not necessarily imply a smaller rate loss, for the commonly used uniform quantizer we prove that the unique loading factor that minimizes the MSE also maximizes the GMI (i.e., achieves zero loading loss). For perfect gain control, we show that the irreducible loss is determined by the product of the normalized MMSE and the SNR, and provide a geometrical interpretation for this result. Performance approximations and per-bit rules in this case are also given. For imperfect gain control, to understand its impact, we characterize the decay of small rate loss in the high-resolution regime, and propose approximations of the achievable rate as a function of the loading factor which are fairly accurate for moderate resolutions. These results provide insight into transceiver design with nonnegligible quantization effect, especially regarding the choice of quantizer resolution and the design of the AGC.
A limitation of this work is that the obtained analytical results are derived under the Gaussian codebook. Like many classical information-theoretic results given by the mutual information, for the GMI, it is also not easy to get closed-form or analytical results without assuming a Gaussian input. For practical systems with finite alphabets, we can expect that our result applies well for Gaussian-like constellations that come from constellation shaping. But it is natural to ask whether our results are still useful in the cases of QAM, phase-shift keying (PSK), and other commonly used inputs, which are all bounded (unlike the unbounded Gaussian input). Since in these cases Proposition 1 does not apply, the GMI evaluation may rely heavily on numerical computation. We left this problem to future study. Nevertheless, the insight gained in this work will be very helpful for the finite-alphabet case.